What is the y-coordinate of the solution to the system of equations?
y = x - 2
y = -x - 4

Answer by Fombitz(32388)

Set the two equations equal to each other.
x-2=-x-4
2x=-2
x=-1
Now use either equation to find the y-coordinate.
y=-1-2
y=-3